OpenStudy (anonymous):

find the numerical derivative of f(x) = 4x-x^2 at the point x=0. use h=0.001

OpenStudy (anonymous):

calculator exercise

OpenStudy (anonymous):

your job is to compute \[\frac{f(0+h)-f(0)}{h}\] but since \(f(0)=0\) and \(0+h=h\) this boils down to computing \[\frac{4h+h^2}{h}\]or since \(h=0.001\) \[\frac{4\times 0.001-(0.001)^2}{0.001}\]

OpenStudy (anonymous):

looks like we are getting pretty damned close to 4 http://www.wolframalpha.com/input/?i= \frac{4\times+0.001-%280.001%29^2}{0.001}

OpenStudy (anonymous):

ooh. okay i understand! your version is alot less complicated than my text book! :)

OpenStudy (anonymous):

lol i think basic, being as how i am

OpenStudy (anonymous):

nahh. you've helped me multiple times! you're great (:

OpenStudy (anonymous):

(blush)